# Early Directional Convergence in Deep Homogeneous Neural Networks for Small Initializations

### v1

@@ A.3 Gronwall’s Inequality
$\frac{d\mathbf{u}}{dt}=\nabla\mathcal{N}_{\mathbf{y},\mathcal{H}}(\mathbf{u})=\mathcal{J}\left(\mathbf{X};\mathbf{u}\right)^{\top}\mathbf{y},\mathbf{u}(0)=\mathbf{w}_{0}.$ (40)
From Lemma 5.2, we know that there are two possibilities. Either $\mathbf{u}(t)$ converges to $\mathbf{0}$, or $\mathbf{u}(t)$ converges in direction to a non-negative KKT point of the constrained NCF. If $\mathbf{u}(t)$ converges to $\mathbf{0}$, then we define $\eta=2.$ Then, for any $\epsilon\in(0,\eta/2)$, we define $B_{\epsilon}=\epsilon$ and choose $T_{\epsilon}$ large enough such that
@@
Else, if $\mathbf{u}(t)$ does not converge to $\mathbf{0}$, then from Lemma B.6, there exists $\tilde{\eta}>0$ and $T$ such that $\|\mathbf{u}(t)\|_{2}\geq\tilde{\eta}$, for all $t\geq T$. We further define $\eta=\min(\tilde{\eta},1)$. Then, from Lemma B.6, for any $\epsilon\in(0,\eta/2)$ there exists $T_{\epsilon}\geq T$ and $B_{\epsilon}$ such that
$\frac{\mathbf{u}(T_{\epsilon})^{\top}{\mathbf{u}^{*}}}{\|\mathbf{u}(T_{\epsilon})\|_{2}}\geq 1-\epsilon,\text{ and }\|\mathbf{u}(t)\|_{2}\leq B_{\epsilon},\forall t\leq T_{\epsilon},$ (42)
where $\mathbf{u}^{*}$ is a non-negative KKT point of
$\max_{\|\mathbf{u}\|_{2}^{2}=1}\mathcal{N}_{\mathbf{y},\mathcal{H}}(\mathbf{u})=\mathcal{H}\left(\mathbf{X};\mathbf{u}\right)^{\top}\mathbf{y}.$
@@
We further define $C_{\epsilon}=\max(\beta\tilde{B}_{\epsilon}^{L},\|\mathbf{y}\|^{2}e^{\|\mathbf{y}\|\beta}\beta\tilde{B}_{\epsilon}^{L})$, and
$\bar{\delta}^{L}=\min\left(\frac{1}{\tilde{B}^{L}_{\epsilon}},\frac{\epsilon}{4\alpha\tilde{B}_{\epsilon}^{L-1}C_{\epsilon}T_{\epsilon}e^{2T_{\epsilon}K_{\epsilon}\|\mathbf{y}\|_{2}}}\right),$ (45)
@@
for square or logistic loss. Then, we can rewrite the above differential equation as
$\dot{\mathbf{w}}=-\mathcal{J}(\mathbf{X};\mathbf{w})^{\top}\mathbf{e},{\mathbf{w}}(0)=\delta\mathbf{w}_{0},$ (48)
where $\mathbf{e}$ is an $n-$dimensional vector such that ${e}_{i}:=\nabla_{\hat{y}}\ell\left(\mathcal{H}(\mathbf{x}_{i};\mathbf{w}),y_{i}\right).$ For square loss, $\nabla_{\hat{y}}\ell(\hat{y},y)=\hat{y}-y$, and for logistic loss, $\nabla_{\hat{y}}\ell(\hat{y},y)=-(2ye^{-\hat{y}y})/(1+e^{-\hat{y}y})$. We define $\xi(t):=\mathbf{e}(t)+\mathbf{y}$, then, for square loss
$\xi(t)=\mathcal{H}(\mathbf{X};\mathbf{w}(t)),$
and for logistic loss
$\xi_{i}(t)=y_{i}\left(\frac{1-e^{-y_{i}\mathcal{H}(\mathbf{x}_{i};\mathbf{w}(t))}}{1+e^{-y_{i}\mathcal{H}(\mathbf{x}_{i};\mathbf{w}(t))}}\right).$
Thus, the dynamics of $\mathbf{w}(t)$ can be written as
$\dot{\mathbf{w}}\in-\mathcal{J}(\mathbf{X};\mathbf{w}(t))^{\top}\mathbf{e}(t)=\mathcal{J}(\mathbf{X};\mathbf{w}(t))^{\top}(\mathbf{y}-\xi(t)).$ (49)
@@
$=\frac{1}{\delta^{L-1}}\mathcal{J}\left(\mathbf{X};\mathbf{w}\left(\frac{t}{\delta^{L-2}}\right)\right)^{\top}\left(\mathbf{y}-\xi\left(\frac{t}{\delta^{L-2}}\right)\right)$
$=\mathcal{J}\left(\mathbf{X};\frac{1}{\delta}\mathbf{w}\left(\frac{t}{\delta^{L-2}}\right)\right)^{\top}\left(\mathbf{y}-\xi\left(\frac{t}{\delta^{L-2}}\right)\right)$
$=\mathcal{J}\left(\mathbf{X};\mathbf{s}(t)\right)^{\top}\left(\mathbf{y}-\xi\left(\frac{t}{\delta^{L-2}}\right)\right),$ (50)
where in third equality, from Lemma B.1, we used $(L-1)$-homogeneity of $\mathcal{J}\left(\mathbf{X};\mathbf{w}\right)$. Using $L$-homogeneity of $\mathcal{H}(\mathbf{X};\mathbf{w})$, we have, for square loss,
$\xi(t/\delta^{L-2})=\mathcal{H}(\mathbf{X};\mathbf{w}(t/\delta^{L-2}))=\delta^{L}\mathcal{H}(\mathbf{X};\mathbf{w}(t/\delta^{L-2})/\delta)=\delta^{L}\mathcal{H}(\mathbf{X};\mathbf{s}(t)),$
and similarly for logistic loss,
$\xi_{i}(t/\delta^{L-2})=y_{i}\left(\frac{1-e^{-y_{i}\delta^{L}\mathcal{H}(\mathbf{x}_{i};\mathbf{s}(t))}}{1+e^{-y_{i}\delta^{L}\mathcal{H}(\mathbf{x}_{i};\mathbf{s}(t))}}\right).$
@@
$\dot{\mathbf{s}}=\mathcal{J}\left(\mathbf{X};\mathbf{s}(t)\right)^{\top}(\mathbf{y}-\tilde{\xi}(t)),\mathbf{s}(0)=\mathbf{w}_{0}.$
@@
$\overline{T}_{\delta}=\inf_{t\in[0,T_{\epsilon}]}\{t:\|\mathbf{s}(t)-\mathbf{u}(t)\|_{2}=\epsilon\}.$
Here, $\overline{T}_{\delta}$ indicates the time when $\|\mathbf{s}(t)-\mathbf{u}(t)\|_{2}$ becomes $\epsilon$ for the first time, and $\delta$ in the subscript indicates that $\overline{T}_{\delta}$ could change with $\delta$. By definition of $\overline{T}_{\delta}$, for all $t\in[0,\overline{T}_{\delta}],$ $\|\mathbf{s}(t)-\mathbf{u}(t)\|_{2}\leq\epsilon$. Now, if we can show $\overline{T}_{\delta}=T_{\epsilon}$, then we would have proved $\|\mathbf{s}(t)-\mathbf{u}(t)\|_{2}\leq\epsilon,$ for all $t\in[0,T_{\epsilon}]$. Since we have assumed $\overline{T}_{\delta}\leq T_{\epsilon}$, we have that $\|\mathbf{u}(t)\|_{2}\leq B_{\epsilon}$ and $\|\mathbf{s}(t)\|_{2}\leq B_{\epsilon}+\epsilon=\tilde{B}_{\epsilon}$, for all $t\in[0,\overline{T}_{\delta}]$. Now, using (44), we have
@@
For square loss and $t\in[0,\overline{T}_{\delta}]$, we have
$\|\tilde{\xi}(t)\|_{2}=\delta^{L}\|\mathcal{H}(\mathbf{X};\mathbf{s}(t))\|_{2}\leq\delta^{L}\beta\|\mathbf{s}(t)\|_{2}^{L}\leq\delta^{L}\beta\tilde{B}_{\epsilon}^{L}.$
Since $\delta\leq 1/\tilde{B}_{\epsilon}$, then for logistic loss and $t\in[0,\overline{T}_{\delta}]$, we have
$|\tilde{\xi}_{i}(t)|=\left|y_{i}\left(\frac{1-e^{-y_{i}\delta^{L}\mathcal{H}(\mathbf{x}_{i};\mathbf{s}(t))}}{1+e^{-y_{i}\delta^{L}\mathcal{H}(\mathbf{x}_{i};\mathbf{s}(t))}}\right)\right|\leq|y_{i}|\left|1-e^{-y_{i}\delta^{L}\mathcal{H}(\mathbf{x}_{i};\mathbf{s}(t))}\right|\leq|y_{i}|e^{|y_{i}|\beta}|y_{i}\delta^{L}\mathcal{H}(\mathbf{x}_{i};\mathbf{s}(t))|\leq|y_{i}|^{2}e^{|y_{i}|\beta}\delta^{L}\beta\tilde{B}_{\epsilon}^{L},$
where the first inequality follows since $|1+e^{x}|\geq 1$. Next, since $\delta\leq 1/\tilde{B}_{\epsilon}$, we have that for all $t\in[0,\overline{T}_{\delta}]$, $\delta^{L}\mathcal{H}(\mathbf{x}_{i};\mathbf{s}(t))\leq\delta^{L}\beta\tilde{B}_{\epsilon}^{L}\leq\beta.$ We obtain the second inequality using the first-order Taylor’s approximation
$\left|1-e^{-y_{i}\delta^{L}\mathcal{H}(\mathbf{x}_{i};\mathbf{s}(t))}\right|\leq e^{|y_{i}|\beta}|y_{i}\delta^{L}\mathcal{H}(\mathbf{x}_{i};\mathbf{s}(t))|.$
For the last inequality we use $\delta^{L}\mathcal{H}(\mathbf{x}_{i};\mathbf{s}(t))\leq\delta^{L}\beta\tilde{B}_{\epsilon}^{L},$ for all $t\in[0,\overline{T}_{\delta}]$. Therefore, $\|\tilde{\xi}(t)\|_{2}\leq\|\mathbf{y}\|^{2}e^{\|\mathbf{y}\|\beta}\delta^{L}\beta\tilde{B}_{\epsilon}^{L}$, for all $t\in[0,\overline{T}_{\delta}].$ Hence, by the definition of $C_{\epsilon}$, for both square loss and logistic loss, we have
@@
$=(\mathbf{s}-\mathbf{u})^{\top}(\mathcal{J}(\mathbf{X};\mathbf{s})-\mathcal{J}(\mathbf{X};\mathbf{u}))^{\top}\mathbf{y}-(\mathbf{s}-\mathbf{u})^{\top}\mathcal{J}(\mathbf{X};\mathbf{s})^{\top}\tilde{\xi}(t)$
$\leq K_{\epsilon}\|\mathbf{y}\|_{2}\|\mathbf{s}-\mathbf{u}\|_{2}^{2}+\alpha\|\mathbf{s}\|_{2}^{L-1}\|\mathbf{s}-\mathbf{u}\|_{2}\|\tilde{\xi}(t)\|_{2}$
$\leq K_{\epsilon}\|\mathbf{y}\|_{2}\|\mathbf{s}-\mathbf{u}\|_{2}^{2}+\alpha\epsilon\tilde{B}_{\epsilon}^{L-1}C_{\epsilon}\delta^{L},$
where the first inequality follows from (51) and (46). For the second inequality we used (52), and $\|\mathbf{s}(t)\|_{2}\leq\tilde{B}_{\epsilon}$, for all $t\in[0,\overline{T}_{\delta}]$. Integrating the above inequality on both sides from 00 to $t_{1}\in[0,\overline{T}_{\delta}]$, we have
@@
Then, using the second Gronwall’s inequality in Lemma A.3, we have
$\|\mathbf{s}(\overline{T}_{\delta})-\mathbf{u}(\overline{T}_{\delta})\|_{2}^{2}\leq 2\alpha\epsilon\tilde{B}_{\epsilon}^{L-1}C_{\epsilon}\delta^{L}\overline{T}_{\delta}e^{2\overline{T}_{\delta}K_{\epsilon}\|\mathbf{y}\|_{2}}.$
For the sake of contradiction assume that $\overline{T}_{\delta}<T_{\epsilon}$. Then, by definition of $\overline{T}_{\delta}$, we know $\|\mathbf{s}(\overline{T}_{\delta})-\mathbf{u}(\overline{T}_{\delta})\|_{2}=\epsilon$. Using the above inequality, and since $\delta<\overline{\delta}$, where $\overline{\delta}$ satisfies (45), we have
$\epsilon^{2}=\|\mathbf{s}(\overline{T}_{\delta})-\mathbf{u}(\overline{T}_{\delta})\|_{2}^{2}\leq 2\alpha\epsilon\tilde{B}_{\epsilon}^{L-1}C_{\epsilon}\delta^{L}T_{\epsilon}e^{2T_{\epsilon}K_{\epsilon}\|\mathbf{y}\|_{2}}\leq\epsilon^{2}/2,$
which leads to a contradiction. Thus, $\overline{T}_{\delta}=T_{\epsilon}$. Since $\|\mathbf{s}(t)-\mathbf{u}(t)\|_{2}\leq\epsilon$, for all $t\in[0,T_{\epsilon}]$, we have
@@
We next prove the second part of the theorem. If $\mathbf{u}(t)$ converges to $\mathbf{0}$, then from (41) we know $\|\mathbf{u}(T_{\epsilon})\|_{2}\leq\epsilon$, which implies
@@
Else, from (42) and (43), we know
$\frac{\mathbf{u}(T_{\epsilon})^{\top}{\mathbf{u}^{*}}}{\|\mathbf{u}(T_{\epsilon})\|_{2}}\geq 1-\epsilon,\text{ and }\|\mathbf{u}(T_{\epsilon})\|_{2}\geq\eta.$ (54)
@@
Next, from (55) we have
@@
$\frac{\mathbf{w}(\overline{T}_{\epsilon})^{\top}{\mathbf{u}}^{*}}{\|\mathbf{w}(\overline{T}_{\epsilon})\|_{2}}$ $=\frac{\mathbf{u}(T_{\epsilon})^{\top}{\mathbf{u}}^{*}+\zeta^{\top}{\mathbf{u}}^{*}}{\|\mathbf{u}(T_{\epsilon})+\zeta\|_{2}}$
$=\left(\frac{\mathbf{u}(T_{\epsilon})^{\top}{\mathbf{u}}^{*}}{\|\mathbf{u}(T_{\epsilon})\|_{2}}\right)\frac{\|\mathbf{u}(T_{\epsilon})\|_{2}}{\|\mathbf{u}(T_{\epsilon})+\zeta\|_{2}}+\frac{\zeta^{\top}{\mathbf{u}}^{*}}{\|\mathbf{u}(T_{\epsilon})+\zeta\|_{2}}$
@@
where in the first inequality we used (54) along with the fact that $\|\zeta\|_{2}\leq\epsilon$ and $\|\mathbf{u}(T_{\epsilon})+\zeta\|_{2}\geq\eta/2.$ The second inequality follows since $\|\zeta\|_{2}\leq\epsilon$ and $\|\mathbf{u}(T_{\epsilon})\|_{2}\geq\eta.$
Appendix C Proofs Omitted from Section 5.2
C.1 Proof of Theorem 5.3

### v2

@@ B.2 Proof of Theorem 1
$\frac{d\mathbf{u}}{dt}=\nabla\mathcal{N}_{-\ell^{\prime}(\mathbf{0},\mathbf{y}),\mathcal{H}}(\mathbf{u})=-\mathcal{J}\left(\mathbf{X};\mathbf{u}\right)^{\top}\ell^{\prime}(\mathbf{0},\mathbf{y}),\mathbf{u}(0)=\mathbf{w}_{0}.$ (42)
From Lemma 2, we know that there are two possibilities. Either $\mathbf{u}(t)$ converges to $\mathbf{0}$, or $\mathbf{u}(t)$ converges in direction to a non-negative KKT point of the constrained NCF.
If $\mathbf{u}(t)$ converges to $\mathbf{0}$, then we define $\eta=2.$ Then, for any $\epsilon\in(0,\eta/2)$, we define $B_{\epsilon}=\epsilon$ and choose $T_{\epsilon}$ large enough such that
@@
Else, if $\mathbf{u}(t)$ does not converge to $\mathbf{0}$, then from Lemma 14, there exists $\tilde{\eta}>0$ and $T$ such that $\|\mathbf{u}(t)\|_{2}\geq\tilde{\eta}$, for all $t\geq T$. We further define $\eta=\min(\tilde{\eta},1)$. Then, from Lemma 14, for any $\epsilon\in(0,\eta/2)$ there exists $T_{\epsilon}\geq T$ and $B_{\epsilon}$ such that
$\frac{\mathbf{u}(T_{\epsilon})^{\top}{\mathbf{u}_{*}}}{\|\mathbf{u}(T_{\epsilon})\|_{2}}\geq 1-\epsilon,\text{ and }\|\mathbf{u}(t)\|_{2}\leq B_{\epsilon},\forall t\leq T_{\epsilon},$ (44)
where $\mathbf{u}_{*}$ is a non-negative KKT point of
$\max_{\|\mathbf{u}\|_{2}^{2}=1}\mathcal{N}_{-\ell^{\prime}(\mathbf{0},\mathbf{y}),\mathcal{H}}(\mathbf{u})=-\mathcal{H}\left(\mathbf{X};\mathbf{u}\right)^{\top}\ell^{\prime}(\mathbf{0},\mathbf{y}).$
@@
Also, since $\nabla_{\hat{y}}\ell(\hat{y},y)$ is locally Lipschitz in $\hat{y}$, there exists $\hat{\beta}$ such that if $\|\mathbf{z}_{1}\|_{2},\|\mathbf{z}_{2}\|_{2}\leq\beta\tilde{B}_{\epsilon}^{L}$, then
$\|\ell^{\prime}\left(\mathbf{z}_{1},\mathbf{y}\right)-\ell^{\prime}\left(\mathbf{z}_{2},\mathbf{y}\right)\|_{2}\leq\hat{\beta}\|\mathbf{z}_{1}-\mathbf{z}_{2}\|_{2},$ (47)
@@
$\beta=\sup\{\|\mathcal{H}(\mathbf{X};\mathbf{w})\|_{2}:\mathbf{w}\in\mathcal{S}^{k-1}\}.$ (48)
We further define $C_{\epsilon}=\beta\hat{\beta}\tilde{B}_{\epsilon}^{L}$, and
$\bar{\delta}^{L}=\min\left(1,\frac{\epsilon}{4\alpha\tilde{B}_{\epsilon}^{L-1}C_{\epsilon}T_{\epsilon}e^{2T_{\epsilon}K_{\epsilon}\|\ell^{\prime}(\mathbf{0},\mathbf{y})\|_{2}}}\right),$ (49)
where recall
$\alpha=\sup\{\|\mathcal{J}(\mathbf{X};\mathbf{w})\|_{2}:\mathbf{w}\in\mathcal{S}^{k-1}\}.$ (50)
@@
then, we can rewrite the above differential equation as
$\dot{\mathbf{w}}=-\sum_{i=1}^{n}\nabla_{\hat{y}}\ell\left(\mathcal{H}(\mathbf{x}_{i};\mathbf{w}),y_{i}\right)\nabla\mathcal{H}(\mathbf{x}_{i};\mathbf{w})=-\mathcal{J}(\mathbf{X};\mathbf{w})^{\top}\ell^{\prime}(\mathcal{H}(\mathbf{X};\mathbf{w}),\mathbf{y}),{\mathbf{w}}(0)=\delta\mathbf{w}_{0}.$ (52)
We define $\xi(t):=\ell^{\prime}(\mathcal{H}(\mathbf{X};\mathbf{w}),\mathbf{y})-\ell^{\prime}(\mathbf{0},\mathbf{y})$, then the dynamics of $\mathbf{w}(t)$ can be written as
$\dot{\mathbf{w}}\in-\mathcal{J}(\mathbf{X};\mathbf{w})^{\top}\ell^{\prime}(\mathcal{H}(\mathbf{X};\mathbf{w}),\mathbf{y})=-\mathcal{J}(\mathbf{X};\mathbf{w})^{\top}(\ell^{\prime}(\mathbf{0},\mathbf{y})+\xi(t)).$ (53)
@@
$=-\frac{1}{\delta^{L-1}}\mathcal{J}\left(\mathbf{X};\mathbf{w}\left(\frac{t}{\delta^{L-2}}\right)\right)^{\top}\left(\ell^{\prime}(\mathbf{0},\mathbf{y})+\xi\left(\frac{t}{\delta^{L-2}}\right)\right)$
$=-\mathcal{J}\left(\mathbf{X};\frac{1}{\delta}\mathbf{w}\left(\frac{t}{\delta^{L-2}}\right)\right)^{\top}\left(\ell^{\prime}(\mathbf{0},\mathbf{y})+\xi\left(\frac{t}{\delta^{L-2}}\right)\right)$
$=-\mathcal{J}\left(\mathbf{X};\mathbf{s}(t)\right)^{\top}\left(\ell^{\prime}(\mathbf{0},\mathbf{y})+\xi\left(\frac{t}{\delta^{L-2}}\right)\right),$ (54)
where in third equality, from Lemma 9, we used $(L-1)$-homogeneity of $\mathcal{J}\left(\mathbf{X};\mathbf{w}\right)$. Using $L$-homogeneity of $\mathcal{H}(\mathbf{X};\mathbf{w})$, we have
$\xi(t/\delta^{L-2})$ $=\ell^{\prime}(\mathcal{H}(\mathbf{X};\mathbf{w}(t/\delta^{L-2})),\mathbf{y})-\ell^{\prime}(\mathbf{0},\mathbf{y})$
$=\ell^{\prime}(\delta^{L}\mathcal{H}(\mathbf{X};\mathbf{w}(t/\delta^{L-2})/\delta),\mathbf{y})-\ell^{\prime}(\mathbf{0},\mathbf{y})$
$=\ell^{\prime}(\delta^{L}\mathcal{H}(\mathbf{X};\mathbf{s}(t)),\mathbf{y})-\ell^{\prime}(\mathbf{0},\mathbf{y}).$
@@
$\dot{\mathbf{s}}=-\mathcal{J}\left(\mathbf{X};\mathbf{s}(t)\right)^{\top}(\ell^{\prime}(\mathbf{0},\mathbf{y})+\tilde{\xi}(t)),\mathbf{s}(0)=\mathbf{w}_{0}.$
@@
$\overline{T}_{\delta}=\inf_{t\geq 0}\{t:\|\mathbf{s}(t)-\mathbf{u}(t)\|_{2}=\epsilon\}.$
Here, $\overline{T}_{\delta}$ indicates the time when $\|\mathbf{s}(t)-\mathbf{u}(t)\|_{2}$ becomes $\epsilon$ for the first time, and $\delta$ in the subscript indicates that $\overline{T}_{\delta}$ could change with $\delta$. By definition of $\overline{T}_{\delta}$, for all $t\in[0,\overline{T}_{\delta}],$ $\|\mathbf{s}(t)-\mathbf{u}(t)\|_{2}\leq\epsilon$. Now, if we can show $\overline{T}_{\delta}>T_{\epsilon}$, then we would have proved $\|\mathbf{s}(t)-\mathbf{u}(t)\|_{2}\leq\epsilon,$ for all $t\in[0,T_{\epsilon}]$.
For the sake of contradiction, suppose $\overline{T}_{\delta}\leq T_{\epsilon}$. This implies that $\|\mathbf{u}(t)\|_{2}\leq B_{\epsilon}$ and $\|\mathbf{s}(t)\|_{2}\leq B_{\epsilon}+\epsilon=\tilde{B}_{\epsilon}$, for all $t\in[0,\overline{T}_{\delta}]$. Now, using eq. 46, we have
@@
Next, since $\delta\leq 1$ and
$\|\mathcal{H}(\mathbf{X};\mathbf{s}(t))\|_{2}\leq\beta\|\mathbf{s}(t)\|_{2}^{L}\leq\beta\tilde{B}_{\epsilon}^{L},\forall t\in[0,\overline{T}_{\delta}],$
using eq. 47, we have
$\|\tilde{\xi}(t)\|_{2}=\|\ell^{\prime}(\delta^{L}\mathcal{H}(\mathbf{X};\mathbf{s}(t)),\mathbf{y})-\ell^{\prime}(\mathbf{0},\mathbf{y})\|_{2}\leq\delta^{L}\hat{\beta}\|\mathcal{H}(\mathbf{X};\mathbf{s}(t))\|_{2}\leq\delta^{L}\hat{\beta}\beta\tilde{B}_{\epsilon}^{L},\forall t\in[0,\overline{T}_{\delta}].$
Hence, by the definition of $C_{\epsilon}$, we have
@@
$=-(\mathbf{s}-\mathbf{u})^{\top}(\mathcal{J}(\mathbf{X};\mathbf{s})-\mathcal{J}(\mathbf{X};\mathbf{u}))^{\top}\ell^{\prime}(\mathbf{0},\mathbf{y})-(\mathbf{s}-\mathbf{u})^{\top}\mathcal{J}(\mathbf{X};\mathbf{s})^{\top}\tilde{\xi}(t)$
$\leq K_{\epsilon}\|\ell^{\prime}(\mathbf{0},\mathbf{y})\|_{2}\|\mathbf{s}-\mathbf{u}\|_{2}^{2}+\alpha\|\mathbf{s}\|_{2}^{L-1}\|\mathbf{s}-\mathbf{u}\|_{2}\|\tilde{\xi}(t)\|_{2}$
$\leq K_{\epsilon}\|\ell^{\prime}(\mathbf{0},\mathbf{y})\|_{2}\|\mathbf{s}-\mathbf{u}\|_{2}^{2}+\alpha\epsilon\tilde{B}_{\epsilon}^{L-1}C_{\epsilon}\delta^{L},$
where the first inequality follows from eq. 55 and eq. 50. For the second inequality we used eq. 56, and $\|\mathbf{s}(t)\|_{2}\leq\tilde{B}_{\epsilon}$, for all $t\in[0,\overline{T}_{\delta}]$. Integrating the above inequality on both sides from 00 to $t_{1}\in[0,\overline{T}_{\delta}]$, we have
@@
Then, using the second Gronwall’s inequality in Lemma 8, we have
$\|\mathbf{s}(\overline{T}_{\delta})-\mathbf{u}(\overline{T}_{\delta})\|_{2}^{2}\leq 2\alpha\epsilon\tilde{B}_{\epsilon}^{L-1}C_{\epsilon}\delta^{L}\overline{T}_{\delta}e^{2\overline{T}_{\delta}K_{\epsilon}\|\ell^{\prime}(\mathbf{0},\mathbf{y})\|_{2}}\leq 2\alpha\epsilon\tilde{B}_{\epsilon}^{L-1}C_{\epsilon}\delta^{L}T_{\epsilon}e^{2T_{\epsilon}K_{\epsilon}\|\ell^{\prime}(\mathbf{0},\mathbf{y})\|_{2}},$
where the second inequality is true since $\overline{T}_{\delta}\leq T_{\epsilon}$. By definition of $\overline{T}_{\delta}$, we know $\|\mathbf{s}(\overline{T}_{\delta})-\mathbf{u}(\overline{T}_{\delta})\|_{2}=\epsilon$. Thus, using the above inequality, and since $\delta<\overline{\delta}$, where $\overline{\delta}$ satisfies eq. 49, we have
$\epsilon^{2}=\|\mathbf{s}(\overline{T}_{\delta})-\mathbf{u}(\overline{T}_{\delta})\|_{2}^{2}\leq 2\alpha\epsilon\tilde{B}_{\epsilon}^{L-1}C_{\epsilon}\delta^{L}T_{\epsilon}e^{2T_{\epsilon}K_{\epsilon}\|\ell^{\prime}(\mathbf{0},\mathbf{y})\|_{2}}\leq\epsilon^{2}/2,$
which leads to a contradiction. Hence, $\overline{T}_{\delta}>T_{\epsilon}$.
Since $\|\mathbf{s}(t)-\mathbf{u}(t)\|_{2}\leq\epsilon$, for all $t\in[0,T_{\epsilon}]$, we have
@@
We next prove the second part of the theorem. If $\mathbf{u}(t)$ converges to $\mathbf{0}$, then from eq. 43 we know $\|\mathbf{u}(T_{\epsilon})\|_{2}\leq\epsilon$, which implies
@@
Else, from eq. 44 and eq. 45, we know
$\frac{\mathbf{u}(T_{\epsilon})^{\top}{\mathbf{u}_{*}}}{\|\mathbf{u}(T_{\epsilon})\|_{2}}\geq 1-\epsilon,\text{ and }\|\mathbf{u}(T_{\epsilon})\|_{2}\geq\eta.$ (58)
@@
Next, from eq. 59 we have
@@
$\frac{\mathbf{w}(\overline{T}_{\epsilon})^{\top}{\mathbf{u}}_{*}}{\|\mathbf{w}(\overline{T}_{\epsilon})\|_{2}}$ $=\frac{\mathbf{u}(T_{\epsilon})^{\top}{\mathbf{u}}_{*}+\zeta^{\top}{\mathbf{u}}_{*}}{\|\mathbf{u}(T_{\epsilon})+\zeta\|_{2}}$
$=\left(\frac{\mathbf{u}(T_{\epsilon})^{\top}{\mathbf{u}}_{*}}{\|\mathbf{u}(T_{\epsilon})\|_{2}}\right)\frac{\|\mathbf{u}(T_{\epsilon})\|_{2}}{\|\mathbf{u}(T_{\epsilon})+\zeta\|_{2}}+\frac{\zeta^{\top}{\mathbf{u}}_{*}}{\|\mathbf{u}(T_{\epsilon})+\zeta\|_{2}}$
@@
where in the first inequality we used eq. 58 along with the fact that $\|\zeta\|_{2}\leq\epsilon$ and $\|\mathbf{u}(T_{\epsilon})+\zeta\|_{2}\geq\eta/2.$ The second inequality follows since $\|\zeta\|_{2}\leq\epsilon$ and $\|\mathbf{u}(T_{\epsilon})\|_{2}\geq\eta.$
Appendix C Proofs omitted from Section 5
We first state some key lemmata which are used to prove Theorem 3 and Theorem 4.
C.1 Key Lemmata
Recall that for any locally Lipschitz continuous function $f:X\to{\mathbb{R}}$, $\nabla f({\mathbf{x}})$ denotes its gradient at ${\mathbf{x}}\in X$, provided $f$ is differentiable at ${\mathbf{x}},$ and its Clarke Subdifferential is denoted by $\partial f({\mathbf{x}})$, which is defined as
$\partial f(\mathbf{x})\coloneqq\text{conv}\left\{\lim_{i\rightarrow\infty}\nabla f(\mathbf{x}_{i}):\lim_{i\rightarrow\infty}\mathbf{x}_{i}=\mathbf{x},\mathbf{x}_{i}\in\Omega\right\},$
where $\Omega$ is any full-measure subset of $X$ such that $f$ is differentiable at each of its points. Note that if $f$ is differentiable everywhere, then $\partial f(\mathbf{x})=\{\nabla f({\mathbf{x}})\}$.
To compute the Clarke subdifferentials of functions which are compositions of non-differentiable functions, we use Clarke’s chain rule of differentiation described in the following lemma
Lemma 15.
(Clarke, 1983, Theorem 2.3.9 ) Let $h_{1},\ldots,h_{n}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ and $g:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be locally Lipschitz functions, and $f(\mathbf{x})=g(h_{1}(\mathbf{x}),\ldots,h_{n}(\mathbf{x}))$, then,
$\partial f(\mathbf{x})\subseteq\text{conv}\left\{\sum_{i=1}^{n}\alpha_{i}\zeta_{i}:\zeta_{i}\in\partial h_{i}(\mathbf{x}),\alpha\in\partial g(h_{1}(\mathbf{x}),\ldots,h_{n}(\mathbf{x}))\right\}.$
To prove our results, we will frequently require the gradients of the NCF with respect to weights of each layer, which are derived next.
Lemma 16.
For $L\geq 2$, let ${\mathcal{H}}$ be an $L$-layer feed-forward neural network whose output is as defined in eq. 9. For all $l\in[L-1]$ and $i\in[n]$, define
$\phi^{0}_{i}={\mathbf{x}}_{i},{\mathbf{h}}_{i}^{1}={\mathbf{W}}_{1}\phi_{i}^{0},$ (60)
$\phi_{i}^{l}=\sigma({\mathbf{h}}_{i}^{l}),{\mathbf{h}}_{i}^{l+1}={\mathbf{W}}_{l+1}\phi_{i}^{l},$ (61)
and let ${\mathbf{A}}_{i}^{l}=\text{diag}\left(\sigma^{\prime}({\mathbf{h}}_{i}^{l})\right)$ [^4] For non-differentiable activation functions such as ReLU, $\sigma^{\prime}$ could be an interval, implying that ${\mathbf{A}}_{i}^{l}$ is a set-valued mapping., where $\sigma^{\prime}(\cdot)$ denotes the Clarke subdifferential of $\sigma(\cdot)$. Also, let
${\mathbf{e}}_{i}^{l}={\mathbf{A}}_{i}^{l}{\mathbf{W}}_{l+1}^{\top}\cdots{\mathbf{A}}_{i}^{L-1}{\mathbf{W}}_{L}^{\top}y_{i}\text{ and }{\mathbf{e}}_{i}^{L}=y_{i},$
for all $l\in[L-1]$ and $i\in[n]$, then
$\partial_{{\mathbf{W}}_{l}}\left(\sum_{i=1}^{n}y_{i}\mathcal{H}({\mathbf{x}}_{i};{\mathbf{W}}_{1},\cdots{\mathbf{W}}_{L})\right)\subseteq\sum_{i=1}^{n}{\mathbf{e}}_{i}^{l}(\phi_{i}^{l-1})^{\top},\text{ for all }l\in[L].$